Mathematics 241 – Multivariable Calculus Quiz 5 – October 19, 2007

Your Name:

A) (5) Let x(t) = (t, t^{2}, t^{3}) be the twisted cubic curve in R^{3}. Set up the integral that would compute the arclength of the segment of this curve for t = 1 to t = 4. (Do not try to compute the numerical value.)

B) (5) Find the curvature function of the curve x(t) = (5 sin(2t), 3, 5 cos(2t)).